# Impact of Sunglasses on One-to-Many Facial Identification Accuracy

### v1

@@ Abstract
One-to-many facial identification is documented to achieve high accuracy in the case where both the probe and the gallery are ‘mugshot quality’ images. However, an increasing number of documented instances of wrongful arrest following one-to-many facial identification have raised questions about its accuracy. Probe images used in one-to-many facial identification are often cropped from frames of surveillance video and deviate from ‘mugshot quality’ in various ways. This paper systematically explores how the accuracy of one-to-many facial identification is degraded by the person in the probe image choosing to wear dark sunglasses. We show that sunglasses degrade accuracy for mugshot-quality images by an amount similar to strong blur or noticeably lower resolution. Further, we demonstrate that the combination of sunglasses with blur or lower resolution results in even more pronounced loss in accuracy. These results have important implications for developing objective criteria to qualify a probe image for the level of accuracy to be expected if it used for one-to-many identification. To ameliorate the accuracy degradation caused by dark sunglasses, we show that it is possible to recover about 38% of the lost accuracy by synthetically adding sunglasses to all the gallery images, without model re-training. We also show that increasing the representation of wearing-sunglasses images in the training set can largely reduce the error rate. The image set assembled for this research will be made available to support replication and further research into this problem.
@@ I Introduction
This paper presents the systematic exploration of the effect of sunglasses on the accuracy of one-to-many facial identification. First, to create a research dataset for this work, we assemble a test set, named “ND-sunglasses”, that contains 16,350 high-quality image pairs (original and with sunglasses), where the only difference in each pair of images is the sunglasses region. Then, we sample a range of degrees of blur and lower resolutions to select ones that have a similar-sized effect on accuracy as sunglasses. Then, we systematically analyze the impact of the combination of sunglasses, blur, and low resolution on accuracy. The results reveal a pronounced accuracy degradation, showing the importance of this study. Next, we introduce a simple but effective approach that can recover up to 37.63% of the accuracy lost, without model re-training, by adding sunglasses to the gallery images. Lastly, we also show that increasing the representation of wearing-sunglasses images in the training set can largely reduce the error rate.
@@
  • Experimental results investigated the impact of adding sunglasses to images in a one-to-many identification scenario. Our findings indicate that probe images with sunglasses exhibited significantly reduced accuracy in one-to-many identification.
  • Analyzing the effect of stacking blur and low resolution on the impact of sunglasses in probe images. In real-world surveillance camera images, a combination of these image quality factors is commonly encountered.
  • Results showing that when a probe image contained sunglasses while the gallery images did not, adding sunglasses to the gallery images can mitigate the accuracy loss due to sunglasses in the probe image.
@@ II Related Work
Figure 2: The pipeline for creating sunglasses-added versions of original probe images. An Android emulator and ManyCam are used to handle the image input and processing steps, the FaceLab app adds sunglasses to the images, and screenshot the FaceLab results to store the images.
The recent NIST report [10] stands as the largest experimental investigation of the accuracy of one-to-many facial identification, given the large number of matchers evaluated and the size of image datasets used. The report emphasizes that demographic differences in the false positive identification rate (FPIR) are significantly larger than those in the false negative identification rate (FNIR), with FPIR differences potentially reaching one or two orders of magnitude (i.e., 10x, 100x). In contrast, FNIR differences are smaller and more specific to the algorithms used. Krishnapriya et al.[19] also examined one-to-many matching accuracy, utilizing the MORPH [26] dataset. Their findings agree with the NIST study, that when the person in the probe has an enrolled image, the one-to-many search is highly likely to return the correct identity. Both [10] and [19] are clear to state that, if the person in the probe image does not have an image enrolled in the gallery, any result returned by one-to-many facial identification must be a false positive identification.
Another NIST report [9] discusses the impressive accuracy achieved when matching mugshot-quality probe images to mugshot-quality enrolled images – “… in galleries containing 12 million individuals, with rank one miss rates approaching 0.1%.” (The term “mugshot quality” refers generally to face images with frontal pose, neutral expression and controlled lighting and background, as also seen with driver’s licence or visa / passport images.) The NIST report notes that the primary datasets comprise good quality images, which is an important factor in the low false positive identification rates achieved by the algorithms. However, this NIST report also notes that, “Recognition in other circumstances, particularly those without a dedicated photographic environment and human or automated quality control checks, will lead to declines in accuracy.”
@@
Bhatta et al. [24] investigate how image quality factors such as blur and reduced resolution affect of one-to-many search using the MORPH dataset. They report that increased blur in the probe image, or reduced resolution of the face in the probe image, can significantly increase the false positive identification rate. They also find a larger accuracy degradation for females than for males.
In addition to image-based factors such as blur and resolution, face images taken from surveillance video may capture a person wearing dark sunglasses. Existing face attribute editing methods could potentially be used to add sunglasses to a face image [13, 20, 11, 28]. However, these methods suffer from unsatisfactory performance on high-resolution images. Moreover, due to the lack of sunglasses labels in popular existing training sets for GANs, adding sunglasses to an image is not (yet) a stable operation with these approaches. Hence, we choose to use a commercial (non-GAN) tool to reliably add realistic-looking sunglasses without changing other regions of the image, and so in an identity-preserving way.
[7] is the most related work, which proposes an occlusion dataset for one-to-many research. Whereas, this dataset has only 161 probe identities and no gender and race labels to address the possible ethical bias, which preventing us from obtaining statistically meaningful results.
This paper extends recent work in exploring the degradation in one-to-many facial identification accuracy caused by wearing sunglasses and providing two possible solutions to mitigate this degradation.
@@ III Dataset
Figure 3: Examples of the original and synthetic sunglasses-added images. Adding synthetic sunglasses alters only the sunglasses region of the original image. Note that the style of the synthetic sunglasses vary between images.
There are specific requirements for a dataset suitable for exploring the effect of sunglasses in probe images on the accuracy of one-to-many identification. One, similar quality images of the same persons, wearing sunglasses in one image and not in another, are required for use as probe images for comparison experiments. Two, additional images of the same persons are required for use as the enrolled gallery images. The additional gallery images of a person should be acquired in a different session than the probe images, to avoid any “same-session effect” in accuracy estimates.
There are existing test sets of face images covering a variety of aspects of facial recognition, including pose [38, 27], age difference [39, 23], size of face in image [5], in-the-wild images and video [15, 18, 31, 22], and illumination and facial hair [35]. None of these focus specifically on the effects of sunglasses. The IJB datasets [18, 31, 22] do contain images of some persons wearing sunglasses. However, the IJB datasets emphasize in-the-wild imagery, which would be appropriate for probe image in our experiments, but not enrolled gallery images. Also, the presence of similar quality images of the same person, both wearing and not wearing sunglasses, would happen only by accident in the IJB datasets. Additionally, NIST has discontinued distribution of the IJB datasets. [^1] https://www.nist.gov/itl/iad/ig/ijb-c-dataset-request-form The AR face image dataset [21] introduced by Martinez in 1998 contains images of persons wearing sunglasses. However, there are only 126 persons in the dataset, the images a grayscale rather than color, and all persons wear the same model/style of sunglasses, images were acquired in two sessions per person, meaning that each person could have only one gallery image in one-to-many identification experiments. In the end, we could not identify an existing dataset suitable for exploring the effects of wearing sunglasses in the context of one-to-many facial identification.
@@
To assemble a suitable dataset for one-to-many identification experiments, we started with the Notre Dame Male/Female Accuracy Dataset (ND-MFAD) [30, 1, 4, 25]. This dataset has several advantages. Images were acquired in a controlled environment over a number of different sessions, so that a gallery of mugshot-quality images can be created. Also, all subjects completed IRB-approved informed consent for their images to be used in research. For our experiments, the images were cropped and aligned using img2pose [2]. The ND-MFAD dataset contains images for 575 subjects who identified as Caucasian female, with 6,669 images total, and for 687 subjects who identified as Caucasian male, totaling 8,419 images.
@@
In this study, we used a commercial app, “FaceLab [^2] https://lyrebirdstudio.net/,” to add sunglasses to each original image. Figure 2 illustrates the pipeline of adding sunglasses to the face images. Our experimental procedure involved processing 15,088 images through FaceLab app using an Android emulator. Occasional technical issues, including emulator crashes and processing failures, resulted in an initial 2% to 3% failure rate. Any failed or incorrect image processing instances were manually processed to have a “sunglasses-added” version for all the original images. Examples of synthetic sunglasses images by FaceLab app are shown in Figure 3.
While the AR dataset is not appropriate for one-to-many identification experiments, it does contain a subset of images featuring subjects wearing real sunglasses. This characteristic allows us to evaluate the identity preservation capability of the FaceLab app when synthetically adding sunglasses to images. To conduct this assessment, we applied synthetic sunglasses to the normal expression images from the AR database using FaceLab. Examples of image comparisons are shown in Figure 4. We then analyzed the cosine similarity distributions, shown in Figure 5, between matching (a) image of a person smiling to image of the person wearing real sunglasses, and (b) same image of a person smiling to image of the person with synthetic sunglasses added. The results show that the distribution of similarity scores with synthetic sunglasses is shifted to higher similarity compared to the distribution with real sunglasses. It is possible that the strong specular highlights in the images with real sunglasses provide an “extra” lowering of the similarity score. In any case, these results provide additional confidence that the operation of adding sunglasses to any image using the FaceLab app is an identity-preserving operation.
The resulting dataset, named “ND-sunglasses”, allows us to conduct experiments where the addition of sunglasses is the sole different between two probe images of a person. To the best of our knowledge, this dataset represents the largest available collection of paired images with and without sunglasses, specifically curated for experiments on one-to-many facial identification.
@@ IV Experiments
The baseline experiment is as follows. For each person in the dataset, their most recent image is taken as their probe image and their older images are their enrolled images for the gallery. For each probe image, we compute its rank-one mated similarity score and rank-one non-mated similarity score. The mated similarity score is the highest similarity between the person’s probe image and one of the same person’s enrolled images. The non-mated similarity score is the highest similarity between the person’s probe image and an enrolled image of a different person. If the non-mated similarity score is higher than the mated similarity score, then this probe image results in a rank-one false positive identification. To consider results across a set of persons, we look at the distributions of mated and non-mated similarity scores. The more separated these distributions are, the generally greater the accuracy of one-to-many identification. To directly focus on the frequency of false positive identifications, we also consider the distribution of (mated – non-mated) scores across a set of probes. Each negative value in this distribution is from a probe that generates a false positive identification. The fraction of such probes gives the false positive identification rate.
@@
The baseline experiment is done with the original mugshot-quality images for probes and enrolled gallery. We then have a series of comparison experiments. Each comparison experiment involves using a modified version of the probe images, matched against the same original mugshot-quality enrolled gallery. For example, the core comparison is made by adding sunglasses to each of the probe images, and observing how the accuracy of one-to-many identification changes compared to the baseline experiment. Making a comparison to the results of the baseline experiment can be done in several ways. One, we can observe the difference in the distributions of the mated scores, non-mated scores, or (mated – non-mated) scores. To give a quantitative estimate to the difference in distributions, we compute the d-prime or the Wasserstein distance [14] between distributions. For example, we can compute the d-prime (or Wasserstein) between the baseline (mated – non-mated) distribution and the (mated – non-mated) distribution for the probe images with sunglasses. A larger d-prime (or Wasserstein) means that sunglasses had a larger effect in degrading accuracy.
The d-prime is commonly used as a quantitative measure of the difference in distributions in biometrics [33, 32, 24]. However, d-prime assumes that the distributions are approximately Gaussian, and the Wasserstein distance (also called “Earth Mover’s Distance) does not. The Wasserstein distance is not currently commonly used in comparing distributions in biometrics, but could be a better choice especially when the distributions are distinctly not Gaussian.
@@
We train two recognition models with the ResNet100 backbone and the Glint360K training set [3]. One is trained with ArcFace loss [6] and one with AdaFace loss [17]. The pattern of accuracy loss due to sunglasses in the probe image is similar between the two matchers. Due to space limits, some ArcFace results are in Supplementary Material.
@@ IV-A Accuracy Degradation by Sunglasses in Probe
The mated and non-mated distributions for the two demographic groups are shown in the supplementary material. The d-prime values derived from these distributions are presented in Table I. In the Adaface results, the d-prime values of mated distributions, 2.6694 for females and 2.8283 for males, suggest a more pronounced effect on males. The d-prime values of non-mated distributions, 0.3003 for females and 0.4357 for males, show that the primary effect is in the mated distributions for probes with sunglasses, both female and male. Due to the quality of the original images and the size of the dataset, the mated and non-mated distributions for probes with sunglasses are still separable, but they have moved closer together. This shows that the presence of sunglasses in a high-quality probe significantly lowers accuracy. And, in a real surveillance scenario, images captured from surveillance video often have lower resolution and/or an increased level of blur. Therefore, we explore the effect of combining blur and lower resolution with sunglasses.
IV-B Effect of Sunglasses + Blur / Resolution

### v2

@@ Abstract
One-to-many facial identification is documented to achieve high accuracy in the case where both the probe and the gallery are ‘mugshot quality’ images. However, an increasing number of documented instances of wrongful arrest following one-to-many facial identification have raised questions about its accuracy. Probe images used in one-to-many facial identification are often cropped from frames of surveillance video and deviate from ‘mugshot quality’ in various ways. This paper systematically explores how the accuracy of one-to-many facial identification is degraded by the person in the probe image choosing to wear dark sunglasses. We show that sunglasses degrade accuracy for mugshot-quality images by an amount similar to strong blur or noticeably lower resolution. Further, we demonstrate that the combination of sunglasses with blur or lower resolution results in even more pronounced loss in accuracy. These results have important implications for developing objective criteria to qualify a probe image for the level of accuracy to be expected if it used for one-to-many identification. To ameliorate the accuracy degradation caused by dark sunglasses, we show that it is possible to recover about 38% of the lost accuracy by synthetically adding sunglasses to all the gallery images, without model re-training. We also show that the frequency of wearing-sunglasses images is very low in existing training sets, and that increasing the representation of wearing-sunglasses images can greatly reduce the error rate. The image set assembled for this research is available at https://cvrl.nd.edu/projects/data/ to support replication and further research.
@@ I Introduction
This paper presents the systematic exploration of the effect of sunglasses on the accuracy of one-to-many facial identification. First, to create a research dataset for this work, we assemble a test set, named “ND-sunglasses”, that contains 15,088 high-quality image pairs (original and with sunglasses), where the only difference in each pair of images is the sunglasses region. Then, we sample a range of degrees of blur and lower resolutions to select ones that have a similar-sized effect on accuracy as sunglasses. Then, we systematically analyze the impact of the combination of sunglasses, blur, and low resolution on accuracy. The results reveal a pronounced accuracy degradation, showing the importance of this study. Next, we introduce a simple but effective approach that can recover up to 38% of the accuracy lost, without model re-training, by adding sunglasses to the gallery images. Lastly, we also show that increasing the representation of wearing-sunglasses images in the training set can greatly reduce the error rate.
@@
  • Experimental results showing that sunglasses in the probe image significantly reduce the accuracy of one-to-many identification.
  • Analysis showing that sunglasses have an additive effect in combination with blur or reduced resolution of the face in the probe image, as may be typical of real-world surveillance camera images.
  • Results showing that when a probe image contains sunglasses while the gallery images do not, adding sunglasses to the gallery images can mitigate the accuracy loss due to sunglasses in the probe.
@@ II Related Work
Figure 2: The pipeline for creating sunglasses-added versions of original images. An Android emulator and ManyCam are used to handle the image input and processing steps, the FaceLab app adds sunglasses, and a screenshot of the FaceLab results is stored.
The NIST report [10] stands as the largest experimental investigation of the accuracy of one-to-many facial identification, given the large number of matchers evaluated and the size of image datasets used. The report emphasizes that demographic differences in the false positive identification rate (FPIR) are significantly larger than those in the false negative identification rate (FNIR), with FPIR differences potentially reaching one or two orders of magnitude (i.e., 10x, 100x). In contrast, FNIR differences are smaller and more specific to the algorithms used. Krishnapriya et al.[19] also examined one-to-many matching accuracy, utilizing the MORPH [26] dataset. Their findings agree with the NIST study, that when the person in the probe has an enrolled image, the one-to-many search is highly likely to return the correct identity. Both [10] and [19] are clear to state that, if the person in the probe image does not have an image enrolled in the gallery, any result returned by one-to-many facial identification must be a false positive identification.
Another NIST report [9] discusses the impressive accuracy achieved when matching mugshot-quality probe images to mugshot-quality enrolled images – “… in galleries containing 12 million individuals, with rank one miss rates approaching 0.1%.” (The term “mugshot quality” refers generally to face images with frontal pose, neutral expression and controlled lighting and background, as also seen with driver’s licence or visa / passport images.) The NIST report notes that the primary datasets comprise good quality images, which is an important factor in the low FPIRs achieved by the algorithms. However, this NIST report also notes that, “Recognition in other circumstances, particularly those without a dedicated photographic environment and human or automated quality control checks, will lead to declines in accuracy.”
@@
Bhatta et al. [24] investigate how image quality factors such as blur and reduced resolution affect one-to-many search using the MORPH dataset. They report that increased blur or reduced resolution of the face in the probe image can significantly increase the FPIR. They also find a larger accuracy degradation for females than for males.
In addition to image-based factors such as blur and resolution, face images taken from surveillance video may capture a person wearing sunglasses. Existing face attribute editing methods could potentially be used to add sunglasses to a face image [13, 20, 11, 28]. However, these methods suffer from unsatisfactory performance on high-resolution images. Moreover, due to the lack of sunglasses labels in popular existing training sets for GANs, adding sunglasses to an image is not (yet) a stable operation with these approaches. Hence, we choose to use a commercial (non-GAN) tool to reliably add realistic-looking sunglasses without changing other regions of the image, so that the addition of sunglasses is identity-preserving.
One related work [7] proposes an occlusion dataset for one-to-many research, but this dataset has only 161 probe identities and lacks gender and race labels to address potential ethical bias, preventing statistically meaningful analysis. Additionally, Wang and Kumar [31] studied how disguises and makeup affect face recognition accuracy, demonstrating significant performance degradation with disguised faces. However, their work focused on general face recognition rather than one-to-many identification.
This paper extends recent work by exploring the degradation in one-to-many facial identification accuracy caused by wearing sunglasses and providing two possible solutions to mitigate this degradation.
@@ III Dataset
Figure 3: Examples of original images and synthetic versions with sunglasses added using FaceLab app. Adding synthetic sunglasses alters only the sunglasses region of the original image. Note that the style of the synthetic sunglasses varies between images.
There are specific requirements for a dataset suitable for exploring the effect of sunglasses in probe images on the accuracy of one-to-many identification. One, similar quality images of the same persons, wearing sunglasses in one image and not in another, are required for use as probe images for comparison. Two, additional images of the same persons are required for use as the enrolled gallery. The gallery images of a person should be acquired in different sessions than the probe images, to avoid any “same-session effect” in accuracy estimates. Three, images should be frontal to reduce the impact of head pose variation, allowing us to focus specifically on how sunglasses affect accuracy.
There are existing test sets of face images covering a variety of aspects of facial recognition, including pose [39, 27], age difference [40, 23], size of face in image [5], in-the-wild images and video [15, 18, 32, 22], and illumination and facial hair [36]. None of these focus specifically on the effects of sunglasses. The IJB datasets [18, 32, 22] contain images of some persons wearing sunglasses. However, the IJB datasets emphasize in-the-wild imagery, which would be appropriate for probe images in our experiments, but not gallery images. Also, the presence of similar quality images of the same person, both wearing and not wearing sunglasses, would happen only by accident in the IJB datasets. Additionally, NIST has discontinued distribution of the IJB datasets. [^1] https://www.nist.gov/itl/iad/ig/ijb-c-dataset-request-form The AR face image dataset [21] introduced by Martinez contains images of persons wearing sunglasses. However, the dataset only contains 126 people with grayscale images rather than color images, all subjects wear the same model/style of sunglasses, and since images were acquired in two sessions per person, each person could have only one gallery image in one-to-many identification experiments In the end, we could not identify an existing dataset suitable for exploring the effects of wearing sunglasses in the context of one-to-many facial identification.
@@
To assemble a suitable dataset for one-to-many identification experiments, we started with the Notre Dame Male/Female Accuracy Dataset (ND-MFAD) [30, 1, 4, 25]. This dataset has several advantages. Images were acquired in a controlled environment over a number of different sessions, so that a gallery of mugshot-quality images can be created. Also, all subjects completed IRB-approved informed consent for their images to be used in research. For our experiments, the images were cropped and aligned using img2pose [2]. The ND-MFAD dataset contains images of 575 subjects who identified as Caucasian female, with a total of 6,669 images, and of 687 subjects who identified as Caucasian male, with a total of 8,419 images.
@@
In this study, we used a commercial app, “FaceLab [^2] https://lyrebirdstudio.net/,” to add sunglasses to each original image. Figure 2 illustrates the pipeline of adding sunglasses to the images. Our experimental procedure involved processing 15,088 images through FaceLab to add sunglasses. Occasional issues, such as emulator crashes and processing failures, resulted in an initial 2% to 3% failure rate. These images were manually processed to have a “sunglasses-added” version for all the original images. Examples of images with synthetic sunglasses added by FaceLab are shown in Figure 3.
While the AR dataset is not appropriate for one-to-many identification experiments, it does contain a subset of images featuring subjects wearing real sunglasses. This characteristic allows us to evaluate the identity preservation capability of the FaceLab app when synthetically adding sunglasses to images. To conduct this assessment, we applied synthetic sunglasses to the normal expression images from the AR database using FaceLab. Examples of image comparisons are shown in Figure 4. We then analyzed the cosine similarity distributions, shown in Figure 5, between matching (a) image of a person smiling to image of the person wearing real sunglasses, and (b) same image of a person smiling to image of the person with synthetic sunglasses added. The results show that the distribution of similarity scores with synthetic sunglasses is shifted to higher similarity compared to the distribution with real sunglasses. It is possible that the strong specular highlights in the images with real sunglasses provide an “extra” lowering of the similarity score. In any case, these results provide additional confidence that the addition of synthetic sunglasses using FaceLab is an identity-preserving operation.
The resulting dataset, named “ND-sunglasses”, allows us to conduct experiments where the addition of sunglasses is the sole difference between two probe images of a person. To the best of our knowledge, this dataset represents the largest available collection of paired images with and without sunglasses, specifically curated for experiments on one-to-many facial identification.
@@ IV Experiments
The baseline experiment is as follows. For each person in the dataset, their most recent image is taken as their probe image and their older images are their enrolled images for the gallery. For each probe image, we compute its rank-one mated similarity score and rank-one non-mated similarity score. The mated similarity score is the highest similarity between the person’s probe image and one of the same person’s enrolled images. The non-mated similarity score is the highest similarity between the person’s probe image and an enrolled image of a different person. If the non-mated similarity score is higher than the mated similarity score, then this probe image results in a rank-one false positive identification (FPI). To consider results across a set of persons, we look at the distributions of mated and non-mated similarity scores. The more separated these distributions are, the generally greater the accuracy of one-to-many identification. To directly focus on the frequency of FPIs, we also consider the distribution of (mated – non-mated) scores across a set of probes. Each negative value in this distribution is from a probe that generates a false positive identification. The fraction of such probes gives the FPIR.
@@
The baseline experiment is done with the original mugshot-quality images for probes and enrolled gallery. We then have a series of comparison experiments. Each comparison experiment involves using a modified version of the probe images, matched against the same original mugshot-quality enrolled gallery. For example, the core comparison is made by adding sunglasses to each of the probe images, and observing how the accuracy of one-to-many identification changes compared to the baseline experiment. Making a comparison to the results of the baseline experiment can be done in several ways. One, we can observe the difference in the distributions of the mated scores, non-mated scores, or (mated – non-mated) scores. To give a quantitative estimate to the difference in distributions, we compute the d-prime or the Wasserstein distance (also called “Earth Mover’s Distance) [14] between distributions. For example, we can compute the d-prime (or Wasserstein) between the baseline (mated – non-mated) distribution and the (mated – non-mated) distribution for the probe images with sunglasses. A larger d-prime (or Wasserstein) means that sunglasses had a larger effect in degrading accuracy.
The d-prime is commonly used as a quantitative measure of the difference in distributions in biometrics [34, 33, 24]. However, d-prime assumes that the distributions are approximately Gaussian, and the Wasserstein distance does not. The Wasserstein distance is not currently commonly used in comparing distributions in biometrics, but could be a better choice especially when the distributions are distinctly not Gaussian.
@@
We use two face recognition models [^3] https://github.com/HaiyuWu/SOTA-Face-Recognition-Train-and-Test pre-trained with the ResNet100 backbone and the Glint360K training set [3]. One is trained with ArcFace loss [6] and one with AdaFace loss [17]. The pattern of accuracy loss due to sunglasses in the probe image is similar between the two matchers. Due to space limits, some ArcFace results are in Supplementary Material.
@@ IV-A Accuracy Degradation by Sunglasses in Probe
The mated and non-mated distributions for the two demographic groups are shown in the supplementary material. The d-prime values derived from these distributions are presented in Table I. In the Adaface results, the d-prime values of mated distributions, 2.6694 for females and 2.8283 for males, suggest a more pronounced effect on males. The d-prime values of non-mated distributions, 0.3003 for females and 0.4357 for males, show that the primary effect is in the mated distributions for probes with sunglasses, both female and male. Due to the quality of the original images and the size of the dataset, the mated and non-mated distributions for probes with sunglasses are still separable, but they have moved closer together. This shows that the presence of sunglasses in a high-quality probe significantly lowers accuracy. And, in a real surveillance scenario, images captured from surveillance video often have lower resolution and/or an increased level of blur. Therefore, we explore the effects of combining blur and lower resolution with sunglasses.
IV-B Effects of Sunglasses + Blur / Resolution